# thanks: These authors contribute equally to this workthanks: These authors contribute equally to this workthanks: These authors contribute equally to this work

### v1

@@ Abstract
where $\mu,\nu=x,y,z$. The protocol generating Eq. (2) imposes specific constraints on the coefficients $\xi_{\mu\nu}$, namely the traceless condition $\sum_{\mu=x,y,z}\xi_{\mu\mu}=0$ and the symmetry condition $\xi_{\mu\nu}=\xi_{\nu\mu}$. An important observation is that if a set $\{\xi_{\mu\nu}\}$ satisfies these constraints, so does the set $\{-\xi_{\mu\nu}\}$. This implies that Floquet engineering always enables reversal of $\hat{H}_{0}$. However, the higher-order terms, denoted by $\hat{H}_{1}$, generally cannot be simultaneously reversed. Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $-\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$. Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics. In this system, $\hat{H}_{1}$ is typically on the order of a few percent relative to $\hat{H}_{0}$ [14].
@@
We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})=(0.025,-0.225,0.2)$, while all other $\xi_{\mu\nu}=0$. For each Hamiltonian $\hat{H}^{(\lambda)}_{0}$, we perform measurements for all nine possible combinations of $\alpha,\gamma\in\{x,y,z\}$ axes, yielding a total of $3\times 9=27$ distinct cases. For each case, we measure three different $\phi$ angles and fit the three curves jointly with a single scramblon ansatz, Eq. (6). We find that all data are excellently captured by the scramblon ansatz Eq. (6), with three representative fits shown in Fig. 2 and the complete set of all data available in the Supplementary Material [45].
@@
Fig. 3 demonstrates how we verify these three constraints for our three Hamiltonians. Panels (a-c) display the values of $a$, $\Delta$, and $\varkappa$ for all nine measurement combinations per Hamiltonian, with dashed lines indicating (i) $\varkappa$ averaged over $\gamma$ and $\alpha$, and (ii) $\alpha$-averaged $a$ and $\Delta$ values that depend only on $\gamma$. Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 13\%$ occasionally, confirming the first two constraints. To test the third constraint, we decompose $M$ into symmetric ($M^{\text{s}}=(M+M^{\text{T}})/2$) and antisymmetric ($M^{\text{a}}=(M-M^{\text{T}})/2$) components. Eigenvalues shown in panels (d-f) reveal that $M^{\text{s}}$ exhibits only one dominant eigenvalue, and all eigenvalues of $M^{\text{a}}$ are negligible, validating the rank-1 symmetric requirement for $M$.
Figure 4: Error Mitigation of the Reversed Quantum Many-Body Dynamics. (a) $F_{\alpha\gamma}(\phi,t)$ and (b) $I_{\alpha\gamma}(t)$ defined in Eq. (8). As an example, we show $\alpha,\gamma=z,y$, $\phi=\pi/64$ and with Hamiltonian $\hat{H}^{(1)}_{0}$. The blue solid line results from the scramblon ansatz by setting $a=0$. The red triangles are error mitigated by $\tilde{F}_{\alpha\gamma}(\phi,t)$, and green circles are raw data without error mitigation. The red dash-dotted line and the green dashed line are obtained from the scramblon ansatze of $\tilde{F}_{\alpha\gamma}(\phi,t)$ and $F_{\alpha\gamma}(\phi,t)$ (without setting $a=0$). The experimental data points in (b) were extracted by fitting the corresponding data for $F_{\alpha\gamma}(\phi,t)$ with a sixth-order polynomial.
The excellent agreement with experimental data strongly validates the scramblon ansatz. Within this framework, the parameter $a$ fully captures imperfections in the time-reversed dynamics, as evidenced by $F_{\alpha\gamma}(0,t)\equiv 1$ when $a=0$. This allows an error mitigation by setting $a=0$ in Eq. (6) while keeping other fitted parameters unchanged (blue curve in Fig. 4a). For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig. 4a). Within the scramblon formalism, this yields:
@@
In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections. We validate this ansatz by fitting it to experimental data from an NMR system with macroscopically randomly interacting spins, where the stringent constraints imposed by scramblon theory are convincingly verified by the fitting outputs. This approach allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics. Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times. Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity.

### v2

@@ Abstract
where $\mu,\nu=x,y,z$. The protocol generating Eq. (2) imposes specific constraints on the coefficients $\xi_{\mu\nu}$, namely the traceless condition $\sum_{\mu=x,y,z}\xi_{\mu\mu}=0$ and the symmetry condition $\xi_{\mu\nu}=\xi_{\nu\mu}$. An important observation is that if a set $\{\xi_{\mu\nu}\}$ satisfies these constraints, so does the set $\{-\xi_{\mu\nu}\}$. This implies that Floquet engineering always enables reversal of $\hat{H}_{0}$. However, the higher-order terms, denoted by $\hat{H}_{1}$, generally cannot be simultaneously reversed. Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $-\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$. Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics. In this system, $\hat{H}_{1}$ and $\delta\hat{H}$ are typically on the order of a few percent relative to $\hat{H}_{0}$ [45, 14].
@@
We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})=(0.2,-0.225,0.025)$, while all other $\xi_{\mu\nu}=0$. For each Hamiltonian $\hat{H}^{(\lambda)}_{0}$, we perform measurements for all nine possible combinations of $\alpha,\gamma\in\{x,y,z\}$ axes, yielding a total of $3\times 9=27$ distinct cases. For each case, we measure three different $\phi$ angles and fit the three curves jointly with a single scramblon ansatz, Eq. (6). We find that all data are excellently captured by the scramblon ansatz Eq. (6), with three representative fits shown in Fig. 2 and the complete set of all data available in the Supplementary Material [45].
@@
Fig. 3 demonstrates how we verify these three constraints for our three Hamiltonians. Panels (a-c) display the values of $a$, $\Delta$, and $\varkappa$ for all nine measurement combinations per Hamiltonian, with dashed lines indicating (i) $\varkappa$ averaged over $\gamma$ and $\alpha$, and (ii) $\alpha$-averaged $a$ and $\Delta$ values that depend only on $\gamma$. Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 16\%$ occasionally, confirming the first two constraints. To test the third constraint, we decompose $M$ into symmetric ($M^{\text{s}}=(M+M^{\text{T}})/2$) and antisymmetric ($M^{\text{a}}=(M-M^{\text{T}})/2$) components. Eigenvalues shown in panels (d-f) reveal that $M^{\text{s}}$ exhibits only one dominant eigenvalue, and all eigenvalues of $M^{\text{a}}$ are negligible, validating the rank-1 symmetric requirement for $M$.
Figure 4: Error Mitigation of the Reversed Quantum Many-Body Dynamics. (a) $F_{\alpha\gamma}(\phi,t)$ and (b) $I_{\alpha\gamma}(t)$ defined in Eq. (8). As an example, we show $\alpha,\gamma=z,y$, $\phi=\pi/64$ and with Hamiltonian $\hat{H}^{(1)}_{0}$. The blue solid line results from the scramblon ansatz by setting $a=0$. The red triangles are error mitigated by $\tilde{F}_{\alpha\gamma}(\phi,t)$, and green circles are raw data without error mitigation. The red dash-dotted line and the green dashed line are obtained from the scramblon ansatze of $\tilde{F}_{\alpha\gamma}(\phi,t)$ and $F_{\alpha\gamma}(\phi,t)$ (without setting $a=0$). The experimental data points in (b) were extracted by fitting the corresponding data for $F_{\alpha\gamma}(\phi,t)$ with a sixth-order polynomial. The errorbars represent the standard deviations.
The excellent agreement with experimental data strongly validates the scramblon ansatz. Within this framework, the parameter $a$ fully captures imperfections in the time-reversed dynamics, as evidenced by $F_{\alpha\gamma}(0,t)\equiv 1$ when $a=0$. This allows an error mitigation by setting $a=0$ in Eq. (6) while keeping other fitted parameters unchanged (blue curve in Fig. 4a). For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig. 4a) [19, 15, 9, 48]. Within the scramblon formalism, this yields:
@@
In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections. We validate this ansatz by fitting it to experimental data from an NMR system with macroscopically randomly interacting spins, where the stringent constraints imposed by scramblon theory are convincingly verified by the fitting outputs. This approach allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics. Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times. Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity [49].